### v1

# Aleph-Alpha-GermanWeb: Improving German-language LLM pre-training with model-based data curation and synthetic data generation

Abstract
Scaling data quantity is essential for large language models (LLMs), yet recent findings show that data quality can significantly boost performance and training efficiency. We introduce a German-language dataset curation pipeline that combines heuristic and model-based filtering techniques with synthetic data generation. We use our pipeline to create Aleph-Alpha-GermanWeb, a large-scale German pre-training dataset which draws from: (1) Common Crawl web data, (2) FineWeb2, and (3) synthetically-generated data conditioned on actual, organic web data. We evaluate our dataset by pre-training both a 1B Llama-style model and an 8B tokenizer-free hierarchical autoregressive transformer (HAT). A comparison on German-language benchmarks, including MMMLU, shows significant performance gains of Aleph-Alpha-GermanWeb over FineWeb2 alone. This advantage holds at the 8B scale even when FineWeb2 is enriched by human-curated high-quality data sources such as Wikipedia. Our findings support the growing body of evidence that model-based data curation and synthetic data generation can significantly enhance LLM pre-training datasets.
1 Introduction
In recent years, ever-larger and more capable large language models (LLMs) have been released. This trend has led to the identification of power-law correlations between the loss value and the number of LLM parameters or pre-training dataset size [20, 47]. As a result, researchers have argued [20] that simply increasing the number of LLM parameters without also increasing the pre-training dataset size yields diminishing returns. Additionally, considering the cost of LLM inference [42] in production use-cases – which can far exceed the model’s initial training cost – practitioners may weigh the cost-benefit trade-off of replacing a larger model with a smaller one to minimise inference costs [22]. To compensate for the reduced model size, a sufficiently-large training data budget would be required [10, 17]. However, this approach poses significant challenges for domains and languages where data is scarce.
Many recent examples in both pre-training [29, 46] and post-training [36, 53] of LLMs indicate that improvements to data quality can make training much more efficient. This is because higher-quality data can reduce the data quantity needed to reach or exceed the same model performance. In fact, conventional data scaling laws, which suggest that a small percentage increase in performance requires increasing the amount of training data by an order of magnitude, appear to be broken when data quality is improved [45]. As a result, improving data quality allows us to train LLMs much more economically. Indeed, a recent 1.7 billion parameter language model pre-trained on 11 trillion tokens achieved new state-of-the-art results for its size by focussing heavily on data quality [3].
Curating datasets to improve their overall quality can be broadly grouped into four categories:
  1. a)
    removing semantic duplication, e.g., otherwise identical documents with slightly different wording, which are typically straightforward to identify and remove [24];
  2. b)
    removing semantic redundancy, e.g., different summaries of the same book, which we expect to maintain some of for the purposes of syntactic variance, but should probably only keep in proportion to the level of importance or complexity of the underlying concept [1];
  3. c)
    removing bad data, e.g., code that doesn’t compile [12] or language which is ungrammatical [38], less our models learn to repeat these mistakes; and
  4. d)
    augmenting or upsampling good data, e.g., generating synthetic data using LLMs [27] or intentionally duplicating high-quality subsets of the data [21].
Many curation methods originated as human-led and heuristic approaches, e.g., preselecting sources with a reputation for high quality data or manual annotation and filtering of datasets [39, 38]. More recently, however, model-based approaches, e.g., training annotation models to classify data into different quality buckets [25] or using LLM-based text quality filtering [46], have been shown to outperform heuristic methods. Additionally augmenting datasets with synthetically-generated data conditioned on high-quality organic data sources not only improves quantity but has been shown to further improve quality [46].
To date, model-based and synthetic approaches exist mostly for English-language text [25, 46]. One of the most commonly-used and largest heuristically-filtered non-English web datasets is FineWeb2 [21], which adapts prior heuristic methods shown to be successful for English web data [39, 38]. Recently, model-based methods have been applied to a subset of languages from FineWeb2 to score them for quality, finding as little as 15% of the original dataset can be used to achieve comparable performance on 1 billion parameter LLMs [31].
Nevertheless, there remains an overall quantity issue; FineWeb2, which includes data from over 1,000 languages, is <20% the size of the English-only FineWeb dataset [38] (8TB vs. 47TB in total disk size). A single language like German constitutes <1.4% of FineWeb2, and model-based approaches only reduce this quantity further [31]. Meanwhile, synthetic approaches to improve the quantity, diversity, and quality of non-English language pre-training data have focussed on machine translation [50], which comes with a host of socio-technical issues [32], and may lack the distinct syntactic or semantic advantages offered by synthetic approaches for LLM pre-training [28]. Here we aim to fill these gaps by constructing a dataset for state-of-the-art German pre-training.
Contributions
In this paper we present a data curation pipeline specifically for German-language LLM pre-training. We use it to derive our own dataset, Aleph-Alpha-GermanWeb, from three sources: (1) Common Crawl [13] web data not included in FineWeb2; (2) FineWeb2; and (3) synthetic data conditioned on actual, organic web data. Our contributions are:
  1. 1.
    To curate the Common Crawl data, we applied a pipeline similar to (but which we show can perform better than) FineWeb2. We then augmented the dataset with synthetically-generated data and applied model-based quality classification methods;
  2. 2.
    We trained two different model architectures, a 1 billion parameter Llama-style model [16] and a tokenizer-free 8 billion parameter model [35], on our dataset and FineWeb2. With these models, we evaluated a range of German language benchmarks, including Multilingual Massive Multitask Language Understanding (MMMLU) [19]. Our evaluations show our dataset outperforms FineWeb2 [21], even when combined with human-curated high-quality data sources such as Wikipedia at the 8 billion parameter scale. Our findings support the growing body of evidence that model-based data curation and synthetic data generation can significantly enhance pre-training datasets; and
  3. 3.
    We make Aleph-Alpha-GermanWeb publicly available to the research community at https://huggingface.co/datasets/Aleph-Alpha/Aleph-Alpha-GermanWeb, in the hope that it will contribute to advancements in German-language LLM pre-training.
2 Methods
Our data curation methodology consists of three approaches: a data filtering pipeline for Common Crawl documents, synthetic document generation, and quality bucketing of all documents.
2.1 Curating Common Crawl
We followed the RefinedWeb pipeline [39], as adapted for German in FineWeb2 [21], with some adjustments. We implemented all steps of the pipeline using NeMo Curator [23]. We summarise the number of web documents our pipeline yielded in Table 1.
CC dump 2024-38 2024-42 2024-46 2024-51 2025-05 2025-08 Total
Original 1,893.59 1,659.31 1,774.51 1,745.08 2,026.56 1,780.59 10,879.64
URL + Lang. 114.14 99.38 107.58 107.29 127.27 108.78 664.43
Content 65.13 55.90 60.75 60.29 72.65 61.20 375.92
Exact Dedup. 45.39 38.05 41.40 47.35 49.65 48.02 269.87
Fuzzy Dedup. 44.53 37.38 40.66 40.07 48.74 42.25 253.64
Global Dedup. Exact 168.19
Global Dedup. Fuzzy 151.61
Table 1: Number of documents (in millions) throughout our curation pipeline across our downloaded Common Crawl data.
Web data corpus.
We downloaded six Common Crawl [13] web data dumps, spanning from September 2024 to February 2025. Each dump captured raw web page data, e.g., HTML, metadata such as URLs, and text extracts.
URL filtering.
We used the URL filter developed in RefinedWeb [39] to filter out: (i) fraudulent and adult websites from a blocklist of 4.6M domains; (ii) additional URLs based on the presence of words associated with fraudulent or adult content; and (iii) content from so-called ‘high-quality’ websites such as Wikipedia, arXiv, etc., which allows these datasets to be mixed in at a later point and in controlled proportions.
Text extraction.
Natural language texts are embedded into raw web data via HTML and other web programming languages. Since we wish to focus on natural language, we extract the natural language from the raw web data using an extraction tool. RefinedWeb used the trafilatura text extractor [5]. However, subsequent work [25] constructing an English-language dataset from Common Crawl data showed that using the resiliparse text extractor [6] performed better on downstream LLM evaluations. We therefore use resiliparse for text extraction.
Language identification.
RefinedWeb [39] used the fastText language classifier of CCNet [51] at the document level. This classifier was trained on character n-grams from Wikipedia and supports 176 natural languages. Using this classifier, we removed documents for which the top language score was not German.
Repetition removal.
Web-scraped documents can include repetitive phrases or words. Heuristic methods for detection and removal of these repetitions were notably developed by Rae et al. [41]. We used these methods with the repetition identification and removal thresholds set in FineWeb2 [21] for German.
Repetition Type Threshold Measurement Unit
Duplicate whole lines $28.2\%$ of lines
Duplicate paragraphs $30.0\%$ of lines
Duplicate paragraphs $20.0\%$ of lines or paragraphs
Lines/paragraphs with many repeated characters $20.0\%$ of lines
Top 2-grams $7.7\%$ of characters
Top 3-grams $10.1\%$ of characters
Top 4-grams $12.3\%$ of characters
All duplicate 5-grams $14.2\%$ of characters
All duplicate 6-grams $12.7\%$ of characters
All duplicate 7-grams $11.5\%$ of characters
All duplicate 8-grams $10.6\%$ of characters
All duplicate 9-grams $9.7\%$ of characters
All duplicate 10-grams $8.8\%$ of characters
Table 2: Repetition removal thresholds for methods introduced in Rae et al. [41], with thresholds set as in FineWeb2 [21] for German.
For example, we discarded the document if it consists of more than 28.2% duplicate lines. This example is shown in the first row of Table 2, where we also list the other repetition types we removed according to different thresholds for document length and document length measurement units.
Document-wise filtering.
We utilized additional document-level heuristics introduced in Rae et al. [41] and adapted in FineWeb2 [21]; namely, we required documents to have:
  1. (i)
    >50 and <100,000 words;
  2. (ii)
    a mean word length of <14 characters;
  3. (iii)
    a symbols-to-words ratio <0.1, where symbols are “#” and “…”;
  4. (iv)
    <90% of lines starting with a bullet point;
  5. (v)
    <30% of lines ending with ellipses;
  6. (vi)
    >77.4% of words containing at least one alphabetic character; and
  7. (vii)
    at least two of the following list of common German words: “der”, “und”, “die”, “in”, “von”, “im”, “den”, “des”, “mit”, “das”, “er”, “dem”, “als”, “wurde”, “für”.
Line-based filtering.
Rae et al. [41] also introduced line-wise corrections to replace or remove unwanted text within documents. We opted instead for a stricter setting by removing documents if:
  1. (i)
    >15% of the document consists of numbers;
  2. (ii)
    >50% of the lines have >50% uppercase characters;
  3. (iii)
    the average words per line is <10 words; or
  4. (iv)
    >40% of paragraphs contain boilerplate strings, e.g., “terms of use”, “privacy policy”, etc..
Exact deduplication.
By first hashing the text of each document, we performed exact deduplication to remove all but one copy of identical documents, i.e., where there are two or more documents whose strings are equal in the dataset, we kept only one of these documents. We performed this deduplication procedure over the entire dataset. This approach is notably unlike FineWeb2 [21], which retained a certain number of duplicate documents according to their duplication rate within the overall dataset, a technique they refer to as ‘rehydration’.
Fuzzy deduplication.
Our document-level fuzzy deduplication methodology closely follows Smith et al. [44], employing a multi-stage process to identify and remove near-duplicates within the corpus. The procedure began with computing MinHash signatures [8] using character-based 5-grams with 23-character sequences, where the number 23 was calculated by assuming an average of 4.5 characters per word. We then used locality sensitive hashing (LSH) [15] to identify candidate duplicates and sort them into 14 buckets, which each had eight unique hashing functions. LSH buckets were then converted to edges for the connected components algorithm. This algorithm identified document clusters with high similarity, producing groups of near-duplicate documents that can then be removed from the corpus. We performed this deduplication procedure over the entire dataset.
2.2 Synthetic data generation
Inspired by similar work for English-language web data [28, 46], we performed synthetic data generation conditioned on actual – what we also refer to as ‘organic’ – data from FineWeb2. We showcase the five prompt templates we used for synthetic data generation in Appendix B, which are designed to generate educational or basic rephrasings, summarizations, question-answer pairs, and lists of factual information. For all synthetic generation, we used the 12 billion parameter Mistral-Nemo-Instruct-2407111https://huggingface.co/mistralai/Mistral-Nemo-Instruct-2407 LLM.
In all cases, if the entire document was longer than a specified character threshold, we segmented it into chunks of characters with lengths equal to or less than the threshold. This was done to prevent the organic document from potentially taking up all remaining context space in the prompt and to improve the quality of the LLM’s responses – prior studies [28, 46] also noted for English-language that if too much organic text is inserted it degrades the synthetic data quality. To avoid splitting the documents mid-word, -sentence, or -paragraph, and thereby potentially giving impractical or nonsensical inputs to the LLM, we segmented the document texts using text-splitter [^2] https://github.com/benbrandt/text-splitter. For a given prompt and document, we synthesized data for all semantic chunks until the organic document’s data was depleted. For each synthesised output, we also performed some post-processing using regular expressions and filtering to remove what we call ‘LLM artifacts’, such as prefixes like “Here’s the rephrased version:” and repetitions of the prompt or input text.
For a large number of FineWeb2 documents, we synthesized data based on one or more of our prompts. This resulted in many documents, especially long ones (which were segmented into more chunks), being represented in multiple ways in the final synthetic dataset. As mentioned in Su, Kong, Lin et al. [46], there is a non-negative (albeit diminishing) performance gain seen with exactly the same data during multi-epoch pre-training [33], therefore it is reasonable to expect that additional ‘synthetic epochs’ conditioned on the same underlying organic data is not harmful up to a similar limit. In the work of Muennighoff et al. [33], this limit was shown to be approximately four epochs. This is why we chose to limit our prompt templates to five, such that there would be no more than five possible synthetic outputs conditioned on the same organic input when the whole dataset was used for training.
2.3 Quality classification
Following prior work on quality classification for English-language web data [46], we categorised documents into five quality buckets – high, medium-high, medium, medium-low, and low – by leveraging the strengths of several imperfect quality classifiers to produce a more accurate outcome. Specifically, we created several individual quality classifiers based on fastText [7] and Bidirectional Encoder Representations from Transformers (BERT) [11] models. In the following, we describe (1) our two grammar classifiers, (2) our two educational quality classifiers, and (3) the way in which we ensembled all classifiers and grouped the data into the five quality buckets according to the classifiers’ scores.
2.3.1 Grammar
We used LanguageTool [^3] https://dev.languagetool.org to annotate a random subset of 400,000 German FineWeb2 documents with the DE_AGREEMENT rule, which identifies text passages with grammatical disagreement. To train our classifiers, we randomly selected 75,000 documents without identified grammar mistakes as high quality examples. As low quality examples, we took 75,000 random documents containing at least one identified grammar error.
We trained a fastText binary classifier on 95% of the data to classify the high and low quality examples, using the remaining 5% for validation. The model reached 63% precision and 63% recall on the validation set. We further trained a BERT-based binary classifier on the same train-validation splits, reaching a precision of 67% and recall of 66% on the validation set.
2.3.2 Educational quality
Inspired by successful demonstrations for English with FineWeb [38] and Nemotron-CC [46], we created educational quality classifiers to curate high-quality German web data by training a classifier on scores given by an LLM-as-a-judge. The intuition behind this is to identify web documents that are more likely to be educational, informative, and useful for model training, as opposed to spam, low-quality, or otherwise unhelpful content.
We first use an LLM-as-a-judge to label a small subset of the data. We then leverage these labelled documents as training data to train more efficient and lightweight fastText and BERT classifiers. This two-stage approach enables us to score a large number of web documents in a computationally-efficient way, while at the same time introducing implicit regularisation to the judging process.
We again made use of Mistral-Nemo-Instruct-2407, this time to annotate a random set of 600,000 documents from German FineWeb2 according to three criteria: (1) content quality, assessing coherence, informativeness and overall quality of the content; (2) language quality, evaluating the use of language, including formality, objectivity, and the presence of errors or slang; and (3) orthography, assessing the correctness of grammar, spelling, and punctuation, including errors such as typos, incorrect verb conjugation, and incorrect declension. We show the prompt used for the LLM-as-a-judge in Appendix A, which provided scores from one to five for each criterion.
For each document, we calculated a combined educational quality score by taking the minimum over the three criteria rated by the LLM-as-a-judge. This is because we observed that the judge would often miss poor quality in one criterium, but rarely in all three. We then used these educational quality scores as the training signal for fastText and BERT quality classification models. We trained one BERT model tasked to predict the educational quality scores given the first 512 tokens of the document’s text. We also trained a binary classifier based on fastText.
For the binary fastText classifier, the low- and high-quality training data subsets consisted of 185,403 documents each. The low quality subset consisted of documents with educational quality scores of one or two, whereas the high quality subset consisted of documents scoring four or five. We used 95% of the data (and the remaining 5% for validation) to train a fastText model to classify between high- and low-quality data. It reached 92% precision and 91.5% recall on the validation set.
To train the BERT classifier, we randomly selected a maximum of 75,000 documents from each class, which we had previously labelled with the LLM-as-a-judge with scores ranging from one to five. The resulting dataset consisted of 75,000 documents for each score above one, and 25,981 documents with scores of one. We used 95% of this dataset for training to predict the one to five scores. The model achieved an overall accuracy of 42% and a macro-average accuracy of 46% when evaluated on the remaining 5% of the data, which served as the validation set.
2.3.3 Ensembling
To split the data into different quality buckets, we use the rules in Table 3 to first allocate overall quality points to each document.
Model Condition Points Initialise all documents 0 BERT (edu. quality) Predicted score = 5 +3 BERT (edu. quality) Predicted score = 4 +2 BERT (edu. quality) Predicted score = 3 +1 fastText (edu. quality) Predicted “high quality” with >99% confidence +3 fastText (edu. quality) Predicted “high quality” with >95% and $\leq$99confidence +2 BERT (grammar) Predicted “high quality” +2 fastText (grammar) Predicted “high quality” with >99% confidence +3 fastText (grammar) Predicted “high quality” with >95% and $\leq$99confidence +2 Document < 100 characters and score > 6 –3
Table 3: Overall quality scoring rules by model, condition, and assigned points.
Using the overall quality scores, we categorised documents into five quality buckets based on the overall quality scores shown in Table 4.
Quality Bucket Score Range Quality Distribution
Filtered CC Synthetic FineWeb2
Docs Bytes Docs Bytes Docs Bytes
High $\geq 10$ 20.7% 10.5% 22.1% 17.9% 16.5% 6.9%
Medium-high $8-9$ 23.7% 18.4% 28.0% 26.4% 19.5% 12.1%
Medium $6-7$ 17.3% 16.7% 21.9% 23.6% 15.4% 11.9%
Medium-low $4-5$ 14.8% 17.2% 14.6% 15.9% 14.8% 14.3%
Low $<4$ 23.4% 37.2% 13.4% 16.3% 33.8% 54.8%
Table 4: Document quality classifications decided by overall quality score.
3 Experiments
We trained and evaluated a one billion parameter (1B) Llama-style transformer model and an eight billion parameter (8B) hierarchical autoregressive transformer (HAT) model on our dataset to assess its quality compared to FineWeb2.
3.1 Evaluations
We measured model performance using the following German-language question-answering tasks.
MMMLU
The Multilingual Massive Multitask Language Understanding dataset [18], is a benchmark designed to evaluate the performance of LLMs across multiple languages and disciplines. It extends the original MMLU benchmark by translating its test set into 14 languages—including German—using professional human translators to ensure accuracy. The dataset encompasses 57 subjects ranging from elementary-level topics to advanced professional fields such as law, physics, history, and computer science. The German portion contains 14,000 samples.
The following three popular LLM benchmarks were translated from English to German [40].
ARC
The AI2 Reasoning Challenge (ARC) [9] is a benchmark dataset developed by the Allen Institute for AI to evaluate the reasoning capabilities of artificial intelligence models. It comprises 7,787 multiple-choice science questions sourced from standardized tests designed for students in grades three through nine. The dataset is divided into two subsets: ARC-Easy and ARC-Challenge; here we evaluate the translation of ARC-Easy.
Figure 1: Average accuracies (MMMLU, ARC-Easy, HellaSwag; single- and five-shot) of 1B Llama-style models trained on $\sim$ 84 billion tokens from different datasets. Our combined dataset, Aleph-Alpha-GermanWeb, and synthetic subset alone, outperforms FineWeb2. Here, 1,000 training steps equates to approximately 2.1 billion tokens. Individual benchmark results are provided in Appendix D.1.
HellaSwag
The HellaSwag benchmark dataset [54] is designed to evaluate the commonsense reasoning abilities of AI models, particularly in the context of sentence completion tasks. It comprises approximately 70,000 multiple-choice questions from diverse sources, including instructional videos and articles from platforms like WikiHow and ActivityNet. Each question presents a context followed by four possible sentence completions, one of which is correct.
TruthfulQA
The TruthfulQA dataset [26] is a benchmark designed to evaluate the truthfulness of language models when generating answers to questions. It comprises 817 questions across 38 categories, including health, law, finance, and politics. The questions are crafted to challenge models with scenarios where humans might hold incorrect beliefs or misconceptions, aiming to assess whether models can avoid generating false answers learned from imitating human texts.
3.2 1B Llama-style model
We pretrained 1B Llama-style models on a total budget of approximately 84 billion tokens from different datasets. Training details can be found in Appendix C.1.
We trained and evaluated our filtered Common Crawl (Filtered CC), synthetic, and an equal mixture of Filtered CC, synthetic, and a subset of FineWeb2 (1-1-1 Mix), and compared these to FineWeb2. The results are shown in Figure 1. We chose to exclude TruthfulQA from the 1B parameter model evaluations, as we found that it did not give a clear signal and suspect that this size of the model may be incapable of the task (cf. §3.3 for results of an 8B model).
3.3 8B HAT model
To evaluate our dataset on a different architecture and a larger scale, we pre-trained 8B hierarchical autoregressive transformer (HAT) models [34] (see Appendix C.2 for training details). With this we leverage recent advances demonstrating that jointly training the tokenizer and model can be a promising alternative [37], aligning with the philosophy of end-to-end learning  [48].
An additional benefit of using an end-to-end, tokenizer-free architecture like HAT is that it debiases comparisons between datasets—comparisons which are a central focus of this paper. Tokenizers have a significant impact on the efficacy of model training [2, 52, 49]. Typically they are trained on large text corpora, which makes models using those tokenizers more effective on datasets that resemble the tokenizer’s training distribution. As a result, using a tokenizer-based model can introduce bias when comparing the effectiveness of different pre-training datasets, favouring those models that align more closely with the tokenizer’s original training data [30]. To avoid this bias, we use a HAT model, which does not rely on any pre-trained tokenizer and instead performs fixed, dataset-agnostic, byte-level string splitting.
Figure 2 compares the performance of these models pre-trained for different datasets and training scenarios. Differently to the 1B Llama-style models, these 8B HAT models operate on words, not tokens, resulting in 1,000 training steps equating to approximately three billion words in this set-up. We confirmed our previous finding of our filtered CC data outperforming FineWeb2 at the 1B scale in Figure 2 (left). We also confirmed that our synthetic data on average outperforms randomly-sampled FineWeb2 data in Figure 2 (centre).
When pre-training LLMs, web data is often mixed from different languages and with data from alternative, curated sources [16], e.g., English and German data might be mixed with additional non-web data from books and encyclopedias. To test the performance of our single highest-performing subset, we compared our synthetic data to mix of 50% FineWeb2 and 50% high-quality curated datasets [^4] In particular: the German National Library, European Union Parliament transcripts, European Patent Office applications, German Text Archive, Project Gutenberg books, JRC-Acquis, NewsAPI, Open Legal Data, Open Web Math, peS2o, PhilPapers, Wikibooks, Wikinews, Wikpedia, Wikisource, and Wikivoyage.. Figure 2 (right) shows our synthetic data outperformed this mixture on our tested benchmarks.
Figure 2: Accuracies of 8B HAT models trained on different datasets. On average, and for all but one individual benchmark comparison, Aleph-Alpha-GermanWeb outperforms FineWeb2. (Left) After 50,000 training steps, equating to $\sim$ 150 billion words. (Centre) After 21,000 training steps, equating to $\sim$ 63 billion words. (Right) After 25,000 training steps on an English-language web-derived dataset, followed by 20,000 training steps on the respective German-language datasets, equating to $\sim$ 75 billion English words followed by $\sim$ 60 billion German words.
4 Discussion and Conclusions
This paper presents a comprehensive data curation pipeline for German-language LLM pre-training, resulting in the creation of our high-quality German dataset, Aleph-Alpha-GermanWeb. By leveraging a combination of Common Crawl web data, FineWeb2, and synthetic data conditioned on organic web data, we demonstrate the effectiveness of our data curation pipeline and synthetic data generation in enhancing pre-training datasets. We trained two different model architectures on Aleph-Alpha-GermanWeb and FineWeb2 on a range of German-language benchmarks. Our results show that our dataset outperforms FineWeb2, even when combined with high-quality human-curated data sources, highlighting the potential of further iterating on existing data curation pipelines and synthetic data generation in advancing German-language LLM pre-training.
We believe there are many improvements and further research directions which will be important in this area. For example, we expect high-quality translation of existing, high-performing English datasets – such as explored in Wang, Lu, et al. [50] – to be a helpful augmentation method. In our manual inspection of machine-translated texts from English to German, however, we found a high number of errors. We therefore caution against training on large amounts of machine-translated data without also rigorously assessing its quality and naturalness.
Relatedly, a more thorough analysis of synthetic data and its apparent quality is needed. Recent work has demonstrated that recursively training models exclusively on synthetic data leads to a phenomenon termed ‘model collapse’ [43], wherein model performance gradually deteriorates over successive generations of training and synthesis. However, mixing synthetic data with organic data appears to avoid this issue [14]. Still, it remains unclear to what degree practitioners should opt for organic rather than synthetic data, or whether synthetic data is merely ‘gaming’ popular evaluation methods by providing data in a more relevant syntax than organic data.
We make Aleph-Alpha-GermanWeb publicly available to the research community, with the hope that it will contribute to further advancements in this field. Our findings support the growing body of evidence on the importance of data curation and synthetic data generation in LLM pre-training, and we believe that our work will have a positive impact on the development of German-language LLMs.
Acknowledgments and Disclosure of Funding
We wish to thank: Johannes Messner, Carina Kauf, and Koen Oostermeijer for additional German language support; members of Aleph Alpha Research who assisted in the development and execution of model training and evaluation; members of the infrastructure support team; the FineWeb2 team for providing an excellent starting point; Common Crawl for freely providing such a valuable resource to the research community; and the NeMo Curator team for releasing their open-source curation tool.
Author contributions
Here we identify the main work-streams associated with this project and list the contributing authors in alphabetical order.
Data preparation: Michael Barlow, Bastian Harren, Volker Stampa, Gregor Ziegltrum
Common Crawl pipeline: Michael Barlow, Thomas F Burns, Bastian Harren, Letitia Parcalabescu,     Volker Stampa
Synthetic data generation: Michael Barlow, Thomas F Burns, Björn Deiseroth
Model-based filtering: Thomas F Burns, Bastian Harren, Letitia Parcalabescu, Stephan Wäldchen,     Gregor Ziegltrum
Quality-based ensembling: Michael Barlow, Letitia Parcalabescu
Experiments and evaluations: Björn Deiseroth
Manuscript writing and editing: Thomas F Burns, Letitia Parcalabescu, Stephan Wäldchen
Research and experimental design: Thomas F Burns, Björn Deiseroth
Appendix A Prompt for LLM-as-a-judge in quality filtering
Educational filter – LLM-as-a-judge prompt SYSTEM: Deine Aufgabe ist es zu bewerten, wie gut die deutsche Sprachqualität eines deutschen Textauszugs ist. Gebe deine Bewertung in folgendem JSON-Format:
     "content_criticism": str (Falls es Schwächen der Antwort im Bezug auf Inhalt gibt, zum Beispiel plötzlicher Themenwechsel, unklarer Sinn, oder fehlende Informationen, nenne und belege sie anhand von Beispielen.),
"content_grade": number (Eine Zahl zwischen 1 und 5, wobei 5 die beste und 1 die schlechteste Bewertung ist:
     5: Sehr kohärent und informativ, sehr hohe Qualität.
     4: Kohärent und informativ.
     3: Kleine Schächen im Inhalt, aber der Sinn kann erfasst werden.
     2: Begrenzte Relevanz oder Genauigkeit, der Sinn kann nicht ganz erfasst werden.
     1: Völlig inkonsistent oder unsinnig.
     "language_criticism": str (Falls es Schwächen der Antwort im Bezug auf Sprache gibt, zum Beispiel Mischung verschiedener Sprachen, Schimpfwörter, unübliche Wortwahl, nenne und belege sie anhand von Beispielen.),
"language_grade": number (Eine Zahl zwischen 1 und 5, wobei 5 die beste und 1 die schlechteste Bewertung ist:
     5: Sehr förmliche, objektive deutsche Sprache.
     4: Wenige, vernachlässigbare Sprachfehler.
     3: Es gibt sprachliche Schwächen, aber der Sinn kann gut erfasst werden.
     2: Stark umgangssprachlich/Sprachmischung.
     1: Grobe Sprachmischung, Schimpfwörter oder schlechte Wortwahl, der Sinn kann nicht ganz erfasst werden.
     "orthography_criticism": str (Falls es Schwächen der Antwort im Bezug auf Grammatik und Rechtschreibung oder Tippfehler gibt, nenne und belege sie anhand von Beispielen. Dazu gehören Worte, die getrennt geschrieben wurden obwohl sie zusammengeschrieben werden sollten, falsche Verbbeugung und falsche Deklination, etc.),
"orthography_grade": number (Eine Zahl zwischen 1 und 5, wobei 5 die beste und 1 die schlechteste Bewertung ist:
     5: Keine Fehler.
     4: wenige unbedeutende Fehler.
     3: einige unbedeutende Fehler.
     2: ein paar grobe Fehler, aber die Antwort ist verständlich.
     1: so viele Fehler, dass die Antwort unverständlich ist.
USER: Textauszug:
{document}
ASSISTANT: Bewertungs-JSON: {
Appendix B Prompts for synthetic data generation
Rephrasing – data generation prompt Geben Sie mir für den folgenden Absatz eine vielfältige Umformulierung in hochwertiger deutscher Sprache, im Stil von Wikipedia. Beginne deine Antwort mit ’Umformulierung:’. Text: {document} Umformulierung:
Summarization – data generation prompt Ihre Aufgabe ist es, den bereitgestellten Text gemäß diesen Anweisungen zu lesen und umzuformulieren: - Versuchen Sie, eine komprimierte, aber genaue und informative Version des Originaltextes zu erstellen, keine vereinfachte Zusammenfassung. - Erfassen und bewahren Sie die entscheidenden Informationen, Schlüsselkonzepte, wichtigen Werte und sachlichen Details im Originaltext auf und machen Sie ihn gleichzeitig lesbarer und zugänglicher. - Behalten Sie Fachbegriffe, Fachvokabular und komplexe Konzepte bei. - Bewahren Sie Beispiele, Erklärungen zu Denkschritten und unterstützende Beweise auf, um die Tiefe und den Kontext des Textes zu erhalten. - Fügen Sie nur Informationen hinzu, die im Originaltext vorhanden sind. Fügen Sie keine neuen oder unbegründeten Behauptungen hinzu. - Schreiben Sie im Klartext Beginne deine Antwort mit ’Umformulierte Version:’. Text: {document} Umformulierte Version:
Rephrasing in Wikipedia style – data generation prompt Ihre Aufgabe ist es, einen neuen Text zu Wissen aus dem bereitgestellten Text zu verfassen, indem Sie diesen Anweisungen folgen: - Schreibe den Text als Passagen mit leicht verständlichen und qualitativ hochwertigen deutsche Sätzen um, wie sie in Lehrbüchern und Wikipedia stehen. - Konzentrieren Sie sich auf inhaltliche Disziplinen wie Geisteswissenschaften, Sozialwissenschaften, Naturwissenschaften, Technik, Ingenieurwesen, Mathematik, Recht und Recht, Wirtschaft, Management, Kunst, Bildung, Agrarwissenschaften, Politik und Geschichte. - Ignorieren Sie Inhalte, die keine nützlichen Fakten oder Kenntnisse enthalten. - Bewahren Sie Beispiele, Erklärungen von Denkprozessen und unterstützende Beweise auf, um die Tiefe und den Kontext des Textes zu erhalten. - Fügen Sie keine Details hinzu oder ändern Sie sie. Wiederholen Sie nur, was bereits im Text steht. - Schreiben Sie im Klartext. - Fügen Sie keine Titel, Untertitel, Notizen oder Kommentare hinzu. Nacheinander wird der Nutzer nun Texte bereitstellen, verfolge dann die Anweisungen und schreibe den neuen Text. Beginne deine Antwort mit ’Neuer Text:’. Text: {document} Neuer Text:
Formulating questions – data generation prompt Lesen Sie den bereitgestellten Text, und stellen Sie allgemeine Fragen. Befolgen Sie die folgenden Anweisungen sorgfältig: 1. Die Fragen sollen auch ohne den Text als Allgemeinwissen beantwortet werden können. 2. Wenn das nicht möglich ist, formuliere zunächst einen kurzen Kontext, oder antworte lediglich ’nicht möglich’. 3. Stellen Sie möglichst diverse Fragen zu unterschiedlichen sachlichen Informationen, wichtigem Wissen oder konkreten Details im Text. 4. Beantworten Sie die Fragen direkt und in kurzer prägnanter Sprache. 5. Stellen Sie Fragen in den folgenden verschiedenen Kategorien: Ja/Nein, Multiple-Choice mit mehreren Optionen zur Auswahl, Vergleichs- und Offener Fragen. 6. Stellen Sie 8 Fragen mit Antworten, zwei zu jeder der beschriebenen Kategorien. 7. Beginne mit ’Fragen-Antwort-Paare:’. Text: {document} Fragen-Antwort-Paare:
Extracting lists – data generation prompt Überprüfen Sie den Text, und extrahieren Sie die wichtigsten Informationen auf Deutsch. Befolgen Sie diese Anweisungen: - Lesen Sie den obigen Text sorgfältig durch und erstellen Sie eine prägnante und organisierte Liste mit sachlichen Informationen, konkreten Details, Schlüsselkonzepten sowie wichtigen Zahlen und Statistiken, die aus dem Text entnommen sind. - Stellen Sie sicher, dass jeder Punkt klar und spezifisch ist und durch den Originaltext gestützt wird. - Stellen Sie sicher, dass der extrahierte Text informationsdicht und leichter zu erlernen ist. - Fügen Sie keine Titel oder Überschriften hinzu. Beginne deine Antwort mit ’Liste:’. Text: {document} Liste:
Appendix C Experiments
C.1 1B Llama-like model training details
We trained a 1 billion parameter transformer model with 16 layers, 2,048 hidden dimensions, 32 attention heads, and 8 key-value heads. The model uses a context length of 4096 tokens and employs modern architecture choices: rotary positional embeddings (RoPE) with a base of 500,000, SwiGLU activation functions with an MLP expansion factor of 4, and RMSNorm for layer normalization.
The model was trained for 40,000 iterations with a global batch size of 512. We used the AdamW optimizer with $\beta_{1}=0.9$, $\beta_{2}=0.95$, $\epsilon=1e-8$, and gradient clipping at 1.0. The learning rate followed a cosine decay schedule with an initial rate of 3e-4, 400 warmup steps, 4,000 cooldown steps, and minimum learning rate of 0.0. We implemented ZeRO optimization for distributed training efficiency. The model uses BFloat16 precision and was trained using a UTF-8 tokenizer with a vocabulary size of 131,072 tokens.
We used the Pharia-1 tokenizer that was trained with the Unigram algorithm using the SentencePiece library[4].
Optim. $\beta_{1}$ $\beta_{2}$ $\epsilon$ $N_{\text{Warmup}}$ $N_{\text{Cooldown}}$ learning rate
AdamW 0.9 0.95 1e-8 1% 10% 3e-4
Table 5: Parameters of the optimizer for the 1B-Llama-model.
$N_{\text{vocab}}$ $n_{\text{Layers}}$ $n_{\text{heads}}$ $d_{\text{model}}$ $d_{\text{MLP}}$ Batch Size Sequence Length
131,072 16 32 2,048 8,192 512 4,096
Table 6: Hyperparameters for the 1B-Llama-model.
C.2 8B HAT model training details
For the second part of our experiments, we trained 8B HAT models that consist of 3 consecutive transformers. In contrast to classical tokenizer-based models, HAT splits words by a rule-based splitter. Words are processed byte-wise by the smaller outer transformers, while the bigger backbone acts as the next-word predictor on a single embedding per word. The first and third models act as encoder and decoder, that transform bytes to word embeddings and vice versa. These models are connected by linear projection layers.
Words are split by special characters and punctuation, and converted into UTF-8 byte sequences.
All three models are trained simultaneously end-to-end, with hyperparameters similar to those in §C.2 and shown in the following tables.
Optim. $\beta_{1}$ $\beta_{2}$ $\epsilon$ $N_{\text{Warmup}}$ $N_{\text{Cooldown}}$ learning rate
AdamW 0.9 0.95 1e-8 1% 10% 1e-4
Table 7: Shared parameters of the optimizer for all transformer modules assembling the 8B-HAT-model.
$N_{\text{vocab}}$ $n_{\text{Layers}}^{\text{Encoder}}$ $n_{\text{Layers}}^{\text{Decoder}}$ $n_{\text{heads}}$ $d_{\text{model}}$ $d_{\text{MLP}}$ Batch Size Sequence Length
256 6 4 6 768 2,047 1,024 variable
Table 8: Hyperparameters for the Encoder/ Decoder Models of 8B-HAT. Almost all parameters are the same for the Encoder and Decoder. Only the number of layers is increased to 6 for the Encoder. As we fixate the number of words per training steps in the Backbone, the Encoder and Decoder will see a variable sequence length per trainingsbatch.
$N_{\text{vocab}}$ $n_{\text{Layers}}$ $n_{\text{heads}}$ $d_{\text{model}}$ $d_{\text{MLP}}$ Batch Size Sequence Length
- 32 32 4,096 11,008 1,024 3,500
Table 9: Hyperparameters for the Backbone Model of 8B-HAT.
Appendix D Extended results
D.1 1B Llama-style model individual benchmarks results
Overall, subsets of Aleph-Alpha-GermanWeb and their combination outperform FineWeb2 across all benchmarks. Interestingly, we find that different subsets have different strengths: our synthetic and high-quality classified data performed particularly well on MMMLU, whereas our filtered CC data performed best for HellaSwag.
Figure 3: Individual benchmark results for 1B Llama-style models trained on $\sim$ 84 billion tokens from different datasets. Here, 1,000 training steps equates to approximately 2.1 billion tokens.